Required information [The following information applies to the questions displayed below.] The following information is avail*Do not consider any underapplied or overapplied overhead. Raw Materials (RM) 47,000 17,000 Indirect materials 189,000 DM useIncome statement (partial) Sales Cost of goods sold Gross profit1. Direct labor usage. 2. Indirect labor usage. 3. Total payroll paid in cash. Prepare journal entries for the above transact1. Incurred other overhead costs (record credit to Other Accounts). 2. Applied overhead to work in process. Prepare journal e
Answer #1
Raw materials (RM) Work in Process (WIP)
RM-April 30 47000 17000 Indirect materials WIP-April 30 9100 389600 Cost of goods manuf.
RM purchases 189000 161000 DM used DM used 161000
DL used 154000
Overhead applied 84700
RM-May 31 58000 WIP-May 31 19200
Finished goods (FG) Factory Overhead
FG-April 30 66000 421200 Cost of goods sold Indirect materials 17000 84700 Overhead applied
Cost of goods manuf. 389600 Indirect labor 46000
Other overhead costs 102000
FG-May 31 34400
Underapplied OH 80300
Income Statement (Partial)
Sales 1300000
Cost of goods sold 421200
Gross Profit 878800
2
For Raw materials:
Debit Credit
Raw materials inventory   189000
       Cash 189000
Work in Process inventory 161000 =47000+189000-17000-58000
    Raw materials inventory   161000
Factory overhead 17000
    Raw materials inventory   17000
3
For Labor:
Debit Credit
Work in Process inventory 154000 =200000-46000
      Factory wages payable 154000
(To record direct labor usage)
Factory overhead 46000
      Factory wages payable 46000
(To record indirect labor usage)
Factory wages payable 200000
       Cash 200000
(To record payroll paid in cash)
4
For overhead:
Debit Credit
Factory overhead 102000
      Other accounts 102000
(To record other overhead)
Work in Process inventory 84700 =(200000-46000)*55%
     Factory overhead 84700
(To record applied overhead)